You are performing a left-tailed t - test with test statistic t = − 0.85 and a sample of size 24, find the p-value to 4 decimal places
Answer #1

Solution:

This is the left tailed test

n = 24

df = n - 1 = 24 - 1 = 23

P-value = P(t < -0.85) = 0.2020